OpenStudy (anonymous):

Convert the polar equation to rectangular form. r=-3sinθ

jimthompson5910 (jim_thompson5910):

r = -3*sin(theta) r*r = r(-3*sin(theta)) r*r = -3r*sin(theta) r*r = -3(r*sin(theta)) r*r = -3y r^2 = -3y x^2 + y^2 = -3y x^2 + y^2 + 3y = 0

OpenStudy (anonymous):

Okay, I understand it up to that point. I know that it forms a circle, I just don't know how to put it in the standard form so I can find the radius.

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

I have the last equation that you typed out, I just need help putting that into x^2+y^2=r^2 so that I can find the radius of the circle that it forms.

jimthompson5910 (jim_thompson5910):

Let's complete the square for the y terms x^2 + y^2 + 3y = 0 x^2 + y^2 + 3y + 9/4 = 0 + 9/4 x^2 + y^2 + 3y + 9/4 = 9/4 x^2 + (y^2 + 3y + 9/4) = 9/4 x^2 + (y + 3/2)^2 = 9/4 x^2 + (y + 3/2)^2 = (3/2)^2 (x-0)^2 + (y - (-3/2))^2 = (3/2)^2 The equation is now in (x-h)^2 + (y-k)^2 = r^2 form where h = 0 k = -3/2 r = 3/2 So the center is (0, -3/2) The radius is 3/2 units

OpenStudy (anonymous):

Can you briefly explain how you got the 9/4? I haven't really gotten the hang of the whole completing the square process. I forgot how to go about that. Everything else, though, I understand. :)

jimthompson5910 (jim_thompson5910):

well I'm focusing on y^2 + 3y and I want to complete the square for that

jimthompson5910 (jim_thompson5910):

so I take half of the 3 to get 3/2 then I square it to get (3/2)^2 = 9/4

jimthompson5910 (jim_thompson5910):

then add this to both sides

jimthompson5910 (jim_thompson5910):

this will give me y^2 + 3y + 9/4 on the left side which will factor to (y + 3/2)^2

OpenStudy (anonymous):

Oh my goodness, thank you! Everything just clicked for me. Thanks!

jimthompson5910 (jim_thompson5910):

yay lol

jimthompson5910 (jim_thompson5910):

you're welcome